# Non-Hermitian Multipole Skin Effects Challenge Localization

### v1

@@ Abstract
We study the effect of quenched disorder on the non-Hermitian skin effect in systems that conserve a U(1) charge and its associated multipole moments. In particular, we generalize the Hatano-Nelson argument for an Anderson transition in a disordered system with non-reciprocal hopping to the interacting case. When only U(1) charge is conserved, we show that there is a transition between a skin effect phase, in which charges are localized at a boundary, and a many-body localized phase, in which the charges are localized at random positions. In periodic boundary conditions, the skin effect gives way to a delocalized phase with a unidirectional current. If additional multipole moments like dipoles are conserved, we show that the non-Hermitian skin effect remains stable for any amount of disorder. As a result, under periodic boundary conditions, the system is always delocalized regardless of disorder strength.
@@
Introduction.—The theory of Anderson localization is a milestone in condensed matter physics that describes how a disordered potential can trap non-interacting quantum particles Anderson (1958). This effect has been generalized to interacting systems—dubbed many-body localization (MBL)—and challenges our understanding of how the laws of statistical mechanics emerge in a closed system Basko et al. (2006); Gornyi et al. (2005); Nandkishore and Huse (2015); Abanin et al. (2019); Sierant et al. (2025). In practice, however, complete isolation is only approximate. Interactions between a quantum system and its environment can induce dissipation, decoherence, and wavefunction collapse, destroying unitarity. Many of the resulting effects can be captured by non-Hermitian Hamiltonians Xiujuan Zhang and Chen (2022); Yuto Ashida and Ueda (2020).
@@
Non-Hermitian systems host a rich variety of behaviors that are not possible in their Hermitian counterparts, ranging from generalized topological phases to novel types of quantum criticality and entanglement transitions Rudner and Levitov (2009); Hu and Hughes (2011); Borgnia et al. (2020); Lee et al. (2019); Okuma et al. (2020); Gong et al. (2018); Yao and Wang (2018); Chen et al. (2020); Bergholtz et al. (2021); Tang et al. (2021); Li et al. (2018); *Li_Yao_19; Skinner et al. (2019); Zabalo et al. (2022); Chan et al. (2019); Turkeshi et al. (2021). One of the most celebrated phenomena is the non-Hermitian skin effect, where an extensive number of single-particle eigenstates are localized at the boundary of a lattice because of non-reciprocal hopping Hatano and Nelson (1996); *Hatano1997; *Hatano1998NonHermitian; Brouwer et al. (1997); Feinberg and Zee (1999). The skin effect emerges from non-Hermitian topology and realizes a contrasting bulk-boundary correspondence to that of Hermitian topological phases Lee et al. (2019); Okuma et al. (2020). Recently, the skin effect was generalized to systems that conserve multipole moments of a scalar U$(1)$ charge Gliozzi et al. (2024). Multipole conserving systems are intrinsically interacting, and the key signature of the multipole skin effect in an $m$-pole conserving system is eigenstates with an extensive $(m+1)$-pole moment. For example, in dipole-conserving systems, the skin effect maximizes the quadrupole moment (see Fig 1(b)).
@@
Here we investigate the impact of disorder on systems that exhibit the non-Hermitian multipole skin effect. First, we generalize the Hatano-Nelson argument to the interacting case and show that, in the presence of U(1) charge conservation, there is a transition between a non-Hermitian skin effect phase and an MBL phase. In the dynamics, this corresponds to a volume-to-area law entanglement phase transition. Next, we consider multipole-conserving systems and demonstrate that, for any amount of disorder, the multipole skin effect remains stable. This also implies that eigenstates are always delocalized for periodic boundary conditions.
@@
For sufficiently strong disorder, the interacting Hermitian system is believed to host an MBL phase Oganesyan and Huse (2007); Pal and Huse (2010); Luitz et al. (2015); Bera et al. (2015); Devakul and Singh (2015); Sierant et al. (2020); Vardhan et al. (2017); De Tomasi et al. (2017); Doggen et al. (2018). The eigenstates in the MBL phase are adiabatically connected to those at $V=0$ by a quasi-local unitary transformation, and therefore maintain a similar spatial structure Imbrie (2016); Ros et al. (2015); Huse et al. (2014). However, these states cannot be expressed in terms of single-particle orbitals as in Eq. (3), so it is convenient to work with conserved operators instead. In the non-interacting case, the local integrals of motion (LIOMs) associated with the right OBC eigenstates of Eq. (6) are given by
@@
These operators reduce to Eq. (7) as $V\rightarrow 0$ and fall off exponentially away from the localization center $j_{\alpha}$, such that $\tilde{\phi}_{\alpha}(\{x_{i}\})\sim e^{-\text{max}(|x_{i}-j_{\alpha}|)/\xi}$ ($\xi$ is not necessarily the same as in the non-interacting case). All LIOMs commute with the total U$(1)$ charge operator $N=\sum_{j}n_{j}$  [^2] Since both the LIOMs and the charge commute with the Hamiltonian, if $[\mathcal{O}_{\alpha},N]\neq 0$, then there would be a degenerate subspace of states, which is almost impossible due to the onsite disorder. See also Ref. Potter and Vasseur, 2016.. Acting with the similarity transformation, we obtain the corresponding non-Hermitian (right-eigenstate) LIOMs:
@@
In any given term of this expansion, the coefficient of the operator scales as $e^{g\max(x_{i})}e^{-\text{max}(|x_{i}-j_{\alpha}|)/\xi}$, so localization occurs when $g<1/\xi,$ and the skin effect occurs when $g>1/\xi$. Thus, the many-body similarity transformation also explains the existence of non-Hermtian MBL, which was observed in Ref. Hamazaki et al., 2019. We confirm the transition numerically by using the dipole moment as an order parameter (see Fig. 3(a)). When the system is periodic, the skin effect is replaced by a persistent current, and the transition is between MBL and delocalized eigenstates. This transition had been observed numerically in PBC Hamazaki et al. (2019); Mák et al. (2024), though its connection to the skin effect and its explanation in terms of LIOMs were not noted.
@@
Heuristically, there is a competition between two tendencies: localizing charges randomly and localizing dipoles on the boundary. However, charges are zero-dimensional objects while dipoles are one-dimensional. The dipole skin effect can thus arise both from fixed-length dipoles moving to the edge and from the boundary dipoles themselves growing in length. The latter process is unique to dipoles, and provides extra pathways to maximize the quadrupole moment. It is this additional localization mechanism that allows the dipole skin effect to dominate over the localization of charge.
@@
We confirm this numerically by plotting the average quadrupole moment of the eigenstates of $H_{\text{dip}}(g)$ in Fig. 4(a): the system develops an extensive quadrupole moment that grows with system size for any $g\neq 0$. In PBC, the skin effect becomes a persistent dipole current that delocalizes the system. To see this, we plot the fraction of complex energies $f_{\text{imag}}$ in Fig. 4(b). In agreement with our analytical argument, the density of complex energies tends to a finite value with increasing system size. We also investigate the out-of-equilibrium dynamics through the imbalance $\mathcal{I}(t)$ starting from the previously defined Néel state, shown in Fig. 4(c). After a transient, the dynamics is frozen for $g=0,$ but mobile for any amount of non-Hermiticity.
@@
The primacy of the dipole skin effect over MBL also extends to the broader family of multipole skin effects Gliozzi et al. (2024). By analogy to the charge and dipole skin effects, the $m$-pole skin effect involves $m$-poles preferentially hopping to one side to generate an $(m+1)$-pole moment. The analogous similarity transformation is generated by $S=e^{f(g)\sum_{j}j^{m+1}n_{j}}$, and Hermitian LIOMs, which can at most be exponentially localized Imbrie (2016); Thomson and Schiró (2018); Chandran et al. (2015); Ros et al. (2015); Huse et al. (2014); Serbyn et al. (2013); Rademaker et al. (2017), are pushed away from their localization centers by the non-Hermiticity. If $m\geq 1$, the multipole skin effect overcomes disorder in OBC, and delocalization ensues in PBC.
@@
Discussion.—We have seen that the fate of localization in disordered, non-Hermitian systems strongly depends on the underlying symmetries. In non-reciprocal systems with a conserved U$(1)$ charge, there is a transition between localization and the non-Hermitian skin effect. Under PBC, this becomes a delocalization transition, as the skin effect gives way to a persistent current. However, when a multipole moment of the charge is also conserved, the skin effect always dominates over localization. This implies that a PBC system with multipole conservation, which imposes kinetic constraints that hinder particle motion Sala et al. (2020); Feldmeier et al. (2020), and strong disorder, which also tends to freeze particles, is nevertheless delocalized as long as some degree of non-reciprocity is present.
@@
In the main text, we considered a dipole-conserving model with strong enough onsite disorder $W=5.5$ such that the Hermitian version, $H_{\text{dip}}(g=0)$ is many-body localized. We showed dynamical evidence for MBL in Fig. 4(c), where the long-time imbalance starting from the Néel state is seen to saturate at a nonzero value when $g=0$. Here we present further spectral evidence that the Hermitian system with $W=5.5$ is indeed many-body localized, focusing on the symmetry sector containing the Néel state.
The first signature we consider is the energy level statistics. Chaotic Hamiltonians exhibit level repulsion, and midspectrum energies are distributed according to the Gaussian Orthogonal Ensemble (GOE). On the other hand, MBL systems do not have revel repulsion, and midspectrum energies have a Poissonian distribution. The proximity to either limit can be quantified by ratio of consecutive energy gaps,

### v2

@@ Abstract
We study the effect of quenched disorder on the non-Hermitian skin effect in systems that conserve a U(1) charge and its associated multipole moments. In particular, we generalize the Hatano-Nelson argument for a localization transition in disordered, non-reciprocal systems to the interacting case. When only U(1) charge is conserved, we show that there is a transition between a skin effect phase, in which charges cluster at a boundary, and a many-body localized phase, in which charges localize at random positions. In the dynamics of entanglement, this coincides with an area to volume law transition. For systems without boundaries, the skin effect becomes a delocalized phase with a unidirectional current. If dipoles or higher multipoles are conserved, we show that the non-Hermitian skin effect remains stable to arbitrary disorder. Counterintuitively, the system is therefore always delocalized under periodic boundary conditions, regardless of disorder strength.
@@
Introduction.—The theory of Anderson localization is a milestone in condensed matter physics that describes how a disordered potential can trap non-interacting quantum particles Anderson (1958). Its generalization to interacting systems—dubbed many-body localization (MBL)—challenges our understanding of how the laws of statistical mechanics emerge in a closed system Basko et al. (2006); Gornyi et al. (2005); Nandkishore and Huse (2015); Abanin et al. (2019); Sierant et al. (2025). In practice, however, complete isolation is only approximate. Interactions between a quantum system and its environment can induce dissipation, decoherence, and wavefunction collapse, destroying unitarity. Many of the resulting effects can be captured by non-Hermitian Hamiltonians Xiujuan Zhang and Chen (2022); Yuto Ashida and Ueda (2020).
@@
Non-Hermitian systems host a variety of behaviors that are not possible in their Hermitian counterparts, ranging from generalized topological phases to novel entanglement transitions and criticality Rudner and Levitov (2009); Hu and Hughes (2011); Borgnia et al. (2020); Lee et al. (2019); Okuma et al. (2020); Manna and Roy (2023); Yao and Wang (2018); Chen et al. (2020); Bergholtz et al. (2021); Tang et al. (2021); Li et al. (2018); *Li_Yao_19; Skinner et al. (2019); Zabalo et al. (2022); Chan et al. (2019); Turkeshi et al. (2021). One of the most celebrated phenomena is the non-Hermitian skin effect, where an extensive number of single-particle eigenstates are localized at the boundary of a lattice because of non-reciprocal hopping Hatano and Nelson (1996); *Hatano1997; *Hatano1998NonHermitian; Brouwer et al. (1997); Feinberg and Zee (1999). Recently, the skin effect was generalized to systems that conserve multipole moments of a scalar U(1) charge Gliozzi et al. (2024). Multipole conserving systems are intrinsically interacting, and the key signature of the multipole skin effect in an $m$-pole conserving system is eigenstates with an extensive $(m+1)$-pole moment. For example, in dipole-conserving systems, the skin effect maximizes the quadrupole moment (see Fig 1(b)).
@@
Here we investigate the impact of disorder on systems that exhibit the non-Hermitian multipole skin effect. First, we generalize the Hatano-Nelson argument to the interacting case and show that, in the presence of U(1) charge conservation, there is a transition between a non-Hermitian skin effect phase and an MBL phase. In the dynamics, this corresponds to a area-to-volume law entanglement phase transition. Next, we consider multipole-conserving systems and demonstrate that, for any amount of disorder, the multipole skin effect remains stable. Remarkably, this implies that under periodic boundary conditions, all eigenstates are delocalized—even in the presence of strong disorder and stringent mobility constraints imposed by multipole conservation.
@@
For sufficiently strong disorder, the interacting Hermitian system is believed to host an MBL phase Oganesyan and Huse (2007); Pal and Huse (2010); Luitz et al. (2015); Bera et al. (2015); Devakul and Singh (2015); Sierant et al. (2020); Vardhan et al. (2017); De Tomasi et al. (2017); Doggen et al. (2018). The eigenstates in the MBL phase are adiabatically connected to those at $V=0$ by a quasi-local unitary transformation, and therefore maintain a similar spatial structure Imbrie (2016); Ros et al. (2015); Huse et al. (2014); Chandran et al. (2015). However, these states cannot be expressed in terms of single-particle orbitals as in Eq. (3), so it is convenient to work with conserved operators instead. In the non-interacting case, the local integrals of motion (LIOMs) associated with the right OBC eigenstates of Eq. (6) are given by
@@
These operators reduce to Eq. (7) as $V\rightarrow 0$ and fall off exponentially away from the localization center $j_{\alpha}$, such that $\tilde{\phi}_{\alpha}(\{x_{i}\})\sim e^{-\text{max}(|x_{i}-j_{\alpha}|)/\xi}$. All LIOMs commute with the total U(1) charge operator $N=\sum_{j}n_{j}$  [^2] Since both the LIOMs and the charge commute with the Hamiltonian, if $[\mathcal{O}_{\alpha},N]\neq 0$, then there would be a degenerate subspace of states, which is almost impossible due to the onsite disorder. See also Ref. Potter and Vasseur, 2016.. Acting with the similarity transformation, we obtain the corresponding non-Hermitian (right-eigenstate) LIOMs:
@@
In any given term of this expansion, the coefficient of the operator scales as $e^{g\max(x_{i})}e^{-\text{max}(|x_{i}-j_{\alpha}|)/\xi}$, so localization occurs when $g<1/\xi,$ and the skin effect occurs when $g>1/\xi$. Thus, the many-body similarity transformation also explains the existence of non-Hermtian MBL, which was observed in Ref. Hamazaki et al., 2019. We confirm the transition numerically by using the dipole moment as an order parameter (see Fig. 3(a)). When the system is periodic, the skin effect is replaced by a persistent current, and the transition is between MBL and delocalized eigenstates. This transition had been observed numerically in PBC Hamazaki et al. (2019); Mák et al. (2024), though its connection to the skin effect and its explanation in terms of LIOMs were not noted. Note also that a version of the similarity transformation argument was previously used to extract the localization length in a Hermitian MBL system Heußen et al. (2021); O’Brien and Refael (2023).
@@
Heuristically, there is a competition between two tendencies: localizing charges randomly and localizing dipoles on the boundary. However, charges are zero-dimensional objects while dipoles are one-dimensional. The dipole skin effect can thus arise both from fixed-length dipoles moving to the edge and from the boundary dipoles themselves growing in length. The latter process is unique to dipoles, and provides extra pathways to maximize the quadrupole moment. It is this additional localization mechanism that allows the dipole skin effect to dominate over the localization of charge  [^7] See Supplemental Material for a case in which dipoles are local, zero-dimensional objects, leading to a transition.
@@
We verify this numerically by plotting the average quadrupole moment of the eigenstates of $H_{\text{dip}}(g)$ in Fig. 4(a): the average quadrupole moment grows with system size for any $g\neq 0$. In PBC, the skin effect becomes a persistent dipole current that delocalizes the system. To see this, we plot the fraction of complex energies $f_{\text{imag}}$ in Fig. 4(b), which tends to a finite value with increasing system size. We also investigate the out-of-equilibrium dynamics through the imbalance $\mathcal{I}(t)$ starting from the previously defined Néel state, shown in Fig. 4(c). After a transient, the dynamics is frozen for $g=0,$ but mobile for any amount of non-Hermiticity.
@@
The primacy of the dipole skin effect over MBL extends to the broader family of multipole skin effects Gliozzi et al. (2024). The $m$-pole skin effect involves $m$-poles preferentially hopping to one side to maximize the $(m+1)$-pole moment. Repeating our previous arguments, the similarity transformation that achieves this is $S\sim e^{g\sum_{j}j^{m+1}n_{j}}$. Hermitian LIOMs, which are at most exponentially localized, are pushed away from their localization centers by the non-Hermiticity. If $m\geq 1$, the multipole skin effect overcomes disorder in OBC, and delocalization ensues in PBC. Beyond multipole conservation, the skin effect can be generalized for any modulated symmetry charge $\mathcal{Q}=\sum_{j}f(j)\,n_{j}$ Sala et al. (2022), where the scaling of $\mathcal{Q}$ determines its localization behavior.
@@
Discussion.—We have seen that the fate of localization in disordered, non-Hermitian systems strongly depends on the underlying symmetries. Non-reciprocal systems that conserve a U(1) charge host a transition between localization and the non-Hermitian skin effect. Under PBC, the skin effect becomes a persistent current and this is a genuine delocalization transition. On the other hand, when a multipole moment of the charge is also conserved, the skin effect always dominates over localization. Consequently, a PBC system with multipole conservation, which imposes kinetic constraints that hinder particle motion Sala et al. (2020); Feldmeier et al. (2020); Moudgalya et al. (2020), and strong disorder, which also tends to freeze particles, is nevertheless delocalized. Surprisingly, non-hermiticity is enough to circumvent both disorder and kinetic constraints and evade localization.
@@
In the main text, we considered a dipole-conserving model with onsite disorder $W=5.5$, such that the Hermitian version, $H_{\text{dip}}(g=0)$, is many-body localized. We showed dynamical evidence for MBL in Fig. 4(c), where the long-time imbalance starting from the Néel state saturates at a nonzero value when $g=0$. Here we present further spectral evidence that the Hermitian system with $W=5.5$ is indeed many-body localized, focusing on the symmetry sector containing the Néel state.
The first signature we consider is the energy level statistics. Chaotic Hamiltonians exhibit level repulsion, and midspectrum energies are distributed according to the Gaussian Orthogonal Ensemble (GOE). On the other hand, MBL systems do not have level repulsion, and midspectrum energies have a Poissonian distribution. The proximity to either limit can be quantified by ratio of consecutive energy gaps,
@@
Lastly, we confirm that the Hermitian MBL phase at $W=5.5$ has a set of exponentially localized conserved operators. These operators correspond to the LIOMs of Eq. 8 in the main text, and their localization in the Hermitian phase is a key property of MBL. We construct these LIOMS numerically by following Ref. Chandran et al. (2015) and taking the long-time average of a local operator:
$\tilde{n}_{i}:=\lim_{T\rightarrow\infty}\frac{1}{T}\int_{0}^{T}n_{i}(t),$ (S4)
where $n_{i}(t)$ denotes the Heisenberg evolution of the local density under the Hamiltonian. After time averaging, only terms diagonal in the energy basis survive:
[…]
Here the sum is taken over the entire eigenstate basis of $H_{\text{dip}}(0)$, and $\tilde{n}_{i}$ is a conserved operator by construction. Furthermore, the support of $\tilde{n}_{i}$ is concentrated near site $i$ in the MBL phase. The spread of this LIOM can be quantified via the two point correlation function,
[…]
where $\mathcal{D}=2^{L}$ is the Hilbert space dimension. This correlator can be interpreted as the effect that a charge at site $i$ has on site $j$ at late times. In an MBL phase, this effect falls off exponentially as $M_{ij}\sim e^{-|i-j|/\xi}$, coinciding with the exponential localization of the LIOM $\tilde{n}_{i}$.
We show the behavior of $M_{ij}$ in both the MBL phase and the clean limit in Fig. S3. The results are averaged over 32 disorder realizations and over different sites in a periodic chain of length $L=18$. For a LIOM defined using the full energy eigenbasis as in Eq. (S5), $M_{ij}\rightarrow 0$ at large $|i-j|$ Chandran et al. (2015). However, we construct the $\tilde{n}_{i}$ by only summing over the eigenstates in the symmetry sector of the Néel state. In other words, we actually calculate
[…]
Because of the symmetry constraints, $M^{\prime}_{ij}$ does not go to zero at large distances, and we must first subtract a background to see the exponential decay. This background is given by
[…]
the long-distance correlator between two operators that are strictly local in the full Hilbert space, but slightly correlated within the symmetry sector. We therefore plot $M_{ij}=M^{\prime}_{ij}-B$, which shows an exponentially localized LIOM in the localized phase. In the clean phase, on the other hand, the putative LIOM $\tilde{n}_{i}$ is no longer local.
Figure S3: Spatial profile of conserved operators $\tilde{n}_{i}$ of $H_{\text{dip}}(0)$ in a periodic chain with $L=18$. In the clean system (blue), these operators are extended, while in the MBL phase at $W=5.5$ (red) they are exponentially localized LIOMs. The results are shown with both (a) linear and (b) logarithmic scales, and are averaged over disorder and lattice sites.